# Coefficients-Preserving Sampling for Reinforcement Learning with Flow Matching

### v1

@@ 1 Introduction
The paradigm of unsupervised pre-training, followed by supervised fine-tuning and reinforcement learning post-training, has become the new standard for training next-generation deep learning models (Achiam et al. (2023); Ouyang et al. (2022)). Inspired by the application of reinforcement learning in Large Language Models (Gao et al. (2023); Rafailov et al. (2023)), RL algorithms have also been adopted in the image and video generation domains (Black et al. (2023); Miao et al. (2024); Wallace et al. (2024); Dong et al. (2023); Yang et al. (2024)). Recently, a series of algorithms have utilized Group Relative Policy Optimization (GRPO) to optimize for specific rewards (Liu et al. (2025); Xue et al. (2025)), achieving impressive results in metrics such as aesthetics (Kirstain et al. (2023); Wu et al. (2023)), instruction following (Hessel et al. (2021)), and image-to-video consistency (Jiang et al. (2024)).
The standard RL loop comprises three stages: sampling, reward and advantage computation, and policy optimization. A crucial requirement of the sampling stage is to generate a group of highly diverse samples for each prompt. To this end, methods such as Flow-GRPO (Liu et al. (2025)) and Dance-GRPO (Xue et al. (2025)) introduce stochasticity by reformulating the deterministic Ordinary Differential Equation (ODE) of the generative process as a Stochastic Differential Equation (SDE). However, we identify that during training, this SDE-based sampling produces outputs always corrupted by conspicuous noise artifacts (see Figure 1). The rewards, which guide the policy updates, are computed from these noisy samples. Consequently, reward models designed to assess aesthetic quality or human preference often assign inaccurate scores and rankings, thereby misleading the learning process.
To resolve this, we thoroughly investigated the SDE sampling mechanism. Our analysis revealed that the SDE formulation injects greater amount of noise than the original ODE. As the original ODE scheduler is retained, this excess noise accumulates, leading to a non-zero final noise level and visibly noisy outputs. Fundamentally, this problem stems from a mismatch between the SDE’s score function term and the noise level introduced by the Wiener process. Inspired by DDIM (Song et al. (2021a)), we reformulated the noise injection method during sampling to ensure that at every timestep, the noise level of the latent variable remains consistent with the scheduler.
@@ 2.1 Alignment For Large Language Models
With the advent of Large Language Models (LLMs)(Brown et al. (2020); Achiam et al. (2023)), Reinforcement Learning (RL) has garnered renewed attention. The Reinforcement Learning from Human Feedback (RLHF) framework(Ouyang et al. (2022); Gao et al. (2023)), for instance, trains a reward model using human preference data, which in turn fine-tunes the LLM to better align with human expectations. As an alternative to computationally intensive policy gradient methods, Direct Preference Optimization (DPO)(Rafailov et al. (2023)) provides a more streamlined approach that directly trains the model on human preference data. More recently, advanced techniques have been applied to enhance multi-step reasoning. For example, OpenAI-o1 utilizes Proximal Policy Optimization (PPO)(Schulman et al. (2017)) and DeepSeek-R1 employs Group Relative Policy Optimization (GRPO)(Shao et al. (2024)), both using verifiable rewards to improve the models’ capacity for extended reasoning via chain-of-thought.
@@ 2.2 Alignment for Diffusion
Analogous to autoregressive LLMs, both Diffusion(Ho et al. (2020); Song et al. (2021a; b)) and Flow Matching(Lipman et al. (2022); Liu et al. (2022)) models usually construct their outputs via a multi-step sampling process. They can therefore be aligned using similar RL-based techniques. As such, diffusion models are compatible with optimization algorithms including DPO(Wallace et al. (2024); Dong et al. (2023); Yang et al. (2024)), PPO-style policy gradients(Black et al. (2023); Miao et al. (2024); Zhao et al. (2025)), and GRPO(Liu et al. (2025); Xue et al. (2025)). However, unlike token-based models that involve discrete selection steps, the absence of quantization in the diffusion process permits an alternative training paradigm: the direct backpropagation of gradients through the full sampling trajectory Xu et al. (2023).
@@ 2.3 Diffusion Sampler
Efficient sampling is a critical research area for diffusion models. Samplers can be broadly categorized by the numerical methods they adapt. Foundational approaches like DDPM (Ho et al. (2020)) established the paradigm but were slow. Denoising Diffusion Implicit Models (Song et al. (2021a)) provided one of the first major speed improvements by formulating a deterministic sampling process. Subsequently, a significant body of work has focused on applying and adapting sophisticated ordinary differential equation (ODE) solvers. For instance, first-order methods like the Euler solver (Song et al. (2021b)) offer speed at the cost of accuracy, while second-order methods like Heun’s method (Karras et al. (2022)) provide a better balance. High-order solvers, such as DPM-Solver (Lu et al. (2022; 2025)), have become popular for their dramatic reduction in required sampling steps. Another effective approach is the predictor-corrector framework, exemplified by UniPC (Zhao et al. (2023)), which unifies these two steps to achieve robust and high-quality generation with minimal function evaluations.
@@ 3 Preliminaries
Flow Matching Assume that $\bm{x}_{0}\sim X_{0}$ is sampled from the data distribution and $\bm{x}_{1}\sim X_{1}$ is a gaussian noise sample, Rectified Flow (Liu et al. (2022)) interpolates noised sample $\bm{x}_{t}$ as,
@@
where $\hat{\bm{v}}_{\theta}(\bm{x}_{t},t)$ is the estimated velocity. The $\hat{\cdot}$ denotes that the value is estimated in the following article.
Flow-GRPO and Dance-GRPO Reinforcement Learning relies on stochastic sampling to generate diverse samples. Flow-GRPO (Liu et al. (2025)) and Dance-GRPO (Xue et al. (2025)) introduce randomness into Flow Matching by converting the deterministic Flow-ODE into Flow-SDE,
@@
$\hat{A}_{t}^{i}=\frac{R(\bm{x}_{0}^{i})-\text{mean}(\{R(\bm{x}^{i}_{0})\}_{i=1}^{G})}{\text{std}(\{R(\bm{x}_{0}^{i})\}_{i=1}^{G})}.$ (4)
Then GRPO (Shao et al. (2024)) optimizes the policy model by maximizing the following objective,
$\mathcal{L}(\theta)=\mathbb{E}_{\bm{c}\sim\mathcal{C},\{\bm{x}^{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|\bm{c})}\frac{1}{G}\sum\limits_{i=1}^{G}\frac{1}{T}\sum\limits_{t=0}^{T-1}(\mathop{min}(r_{t}^{i}(\theta)\hat{A}_{t}^{i},\text{clip}(r_{t}^{i}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}^{i}))),$ (5)
where $r_{t}^{i}(\theta)=\frac{p_{\theta}(\bm{x}^{i}_{t-1}|\bm{x}_{t}^{i})}{p_{\theta_{\text{old}}}(\bm{x}^{i}_{t-1}|\bm{x}_{t}^{i})}$. Here we omit the KL loss term due to space limit.
@@
where $\epsilon_{t}\sim{\mathcal{N}}({\bm{0}},{\bm{I}})$ is standard Gaussian noise independent of $\epsilon_{\theta}^{(t)}(\bm{x}_{t})$. For other notations, please refer to the DDIM paper (Song et al. (2021a)). When $\sigma_{t}=\sqrt{(1-\alpha_{t-1})/(1-\alpha_{t})}\sqrt{1-\alpha_{t}/\alpha_{t-1}}$, the forward process becomes Markovian, and the generative process becomes a DDPM. When $\sigma_{t}=0$, the resulting model becomes an implicit probabilistic model (DDIM). For other $\sigma_{t}$, we call it DDIM with stochasticity.
@@ 4.1 Coefficients-Preserving Sampling
This rearrangement reveals that a standard ODE step is equivalent to re-interpolating the predicted sample and noise using the new timestep $t-\Delta t$, preserving the linear interpolation structure defined in Equation 1. Referring to Equation 6, we can rewrite the sampling function as,
@@
    The total noise level, defined as the standard deviation of a single multivariate noise or the sum of multiple independent noises, must align with the scheduler for all timesteps.
DDIM sampling is Coefficients-Preserving: From Equation 6, there are two independent noise terms, whose coefficients are $\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}$ and $\sigma_{t}$, so the final noise level is $\sqrt{1-\alpha_{t-1}}$. The sample coefficient is $\sqrt{\alpha_{t-1}}$, so the squared sum of the two coefficients is 1. These two coefficients exactly match the DDIM schedule, whatever $\sigma_{t}$ is. Thus, we say the sampling procedure of DDIM is Coefficients-Preserving Sampling.
@@ 4.2 Flow-SDE is not Coefficients-Preserving Sampling
Recall Equation 3 and rewrite it into the form of Equation 6,
@@
$=\left(1-(t-\Delta t)\right)\hat{\bm{x}}_{0}+(t-\Delta t-\frac{\sigma_{t}^{2}\Delta t}{2t})\hat{\bm{x}}_{1}+\sigma_{t}\sqrt{\Delta t}\bm{\epsilon}$ (9)
@@
Figure 2: The ideal and SDE noise level for Flow-GRPO and Dance-GRPO with $1000$, $16$, and 44 sampling steps. Except for the numerical problem around $t=0$ and $t=1$, the error of the noise level increases as the sampling step decreases.
In Figure 2, we plot the noise level for both Flow-GRPO and Dance-GRPO. As we can see, the noise level mismatch problem is severe for both of them. Moreover, because of the $\frac{\sigma_{t}^{2}\Delta t}{2t}$ term in Equation 9, the error around $t=0$ is large for Dance-GRPO ($\sigma_{t}=\eta$). For Flow-GRPO, $\sigma_{t}=\eta\sqrt{\frac{t}{1-t}}$, the noise level is inaccurate around $t=1$. The problem becomes even worse when the sampling step is low, e.g. $4$ steps for FLUX.1-schnell.
@@ 4.3 Our Solution
The main problem of Flow-SDE is that the reduced noise level $\frac{\sigma_{t}^{2}\Delta t}{2t}$ cannot match the newly added noise level $\sigma_{t}\sqrt{\Delta t}$. Noticing that DDPM also injects noise into the DDIM sampling procedure while preserving the noise level (Figure 3.b), we consider referring to DDIM sampling to solve the problem. Assume that the newly added noise has a variance of $\sigma_{t}^{2}$, the coefficient of predicted noise should be $\sqrt{(t-\Delta t)^{2}-\sigma_{t}^{2}}$ to meet the requirement of the second condition of CPS. In this way, the sampling formulation is,
@@
For the noise level $\sigma_{t}$, the maximum value is $t-\Delta t$, or the sqrtsqrt term would have a negative radicand. To avoid the negative radicand, we propose to set $\sigma_{t}=(t-\Delta t)\sin(\frac{\eta\pi}{2})$. Then the sampling formulation becomes,
@@
To train with GRPO, we also need $p_{\theta}(\bm{x}^{i}_{t-1}|\bm{x}_{t}^{i})$, which is defined as (Liu et al. (2025)),
$\log p_{\theta}(\bm{x}^{i}_{t-1}|\bm{x}_{t}^{i})=\frac{\|\bm{x}_{t-\Delta t}-\mu_{\theta}(\bm{x}_{t},t)\|^{2}}{2\sigma_{t}^{2}}-\log\sigma_{t}-\log\sqrt{2\pi},$ (13)
where $\mu_{\theta}(\bm{x}_{t},t)=\left(1-(t-\Delta t)\right)\hat{\bm{x}}_{0}+(t-\Delta t)\cos(\frac{\eta\pi}{2})\hat{\bm{x}}_{1}$. For each step, the $\log\sigma_{t}-\log\sqrt{2\pi}$ is a constant value that cancels out in $r_{t}^{i}(\theta)=\frac{p_{\theta}(\bm{x}^{i}_{t-1}|\bm{x}_{t}^{i})}{p_{\theta_{\text{old}}}(\bm{x}^{i}_{t-1}|\bm{x}_{t}^{i})}$. Moreover, we removed the $\sigma_{t}$ in the denominator to avoid division by zero in the last timestep. Thus, our definition of log-probability is as simple as,
$\log p_{\theta}(\bm{x}^{i}_{t-1}|\bm{x}_{t}^{i})=\|\bm{x}_{t-1}-\mu_{\theta}(\bm{x}_{t},t)\|^{2}.$ (14)
Figure 3: (a): DDIM deterministic sampling process. Note that $\epsilon$ is a random Gaussian noise, which is almost orthogonal to the sample $\bm{x}_{0}$. Since $\sqrt{\alpha_{t}}^{2}+\sqrt{\beta_{t}}^{2}=1$, the trajectory is part of a quarter-circle at each step. (b): DDIM sampling process with stochasticity (Equation 6). $\epsilon_{t}$ is also a random Gaussian noise, which is almost orthogonal to $\bm{x}_{0}$ and $\epsilon_{\theta}$. (c): Flow matching ODE Sampler. The trajectory is a straight line at each step. (d): Our proposed Coefficients-Preserving Sampling (Equation 12). This figure is inspired by Wang (2024).

### v4

@@ 1 Introduction
The paradigm of unsupervised pre-training, followed by supervised fine-tuning and reinforcement learning post-training, has become the new standard for training next-generation deep learning models (achiam2023gpt; ouyang2022training). Inspired by the application of reinforcement learning in Large Language Models (gao2023scaling; rafailov2023direct), RL algorithms have also been adopted in the image and video generation domains (black2023training; miao2024training; wallace2024diffusion; dong2023raft; yang2024using). Recently, a series of algorithms have utilized Group Relative Policy Optimization (GRPO) to optimize for specific rewards (liu2025flow; xue2025dancegrpo), achieving impressive results in metrics such as aesthetics (kirstain2023pick; wu2023human), instruction following (hessel2021clipscore), and image-to-video consistency (jiang2024anisora).
The standard RL loop comprises three stages: sampling, reward and advantage computation, and policy optimization. A crucial requirement of the sampling stage is to generate a group of highly diverse samples for each prompt. To this end, methods such as Flow-GRPO (liu2025flow) and Dance-GRPO (xue2025dancegrpo) introduce stochasticity by reformulating the deterministic Ordinary Differential Equation (ODE) of the generative process as a Stochastic Differential Equation (SDE). However, we identify that during training, this SDE-based sampling produces outputs always corrupted by conspicuous noise artifacts (see Figure 1). The rewards, which guide the policy updates, are computed from these noisy samples. Consequently, reward models designed to assess aesthetic quality or human preference often assign inaccurate scores and rankings, thereby misleading the learning process.
To resolve this, we thoroughly investigated the Flow-SDE sampling mechanism. Our analysis revealed that the Flow-SDE formulation injects a greater amount of noise than the original ODE. As the original ODE scheduler is retained, this excess noise accumulates, leading to a non-zero final noise level and visibly noisy outputs. Fundamentally, this problem stems from a mismatch between the SDE’s score function term and the noise level introduced by the Wiener process. Inspired by DDIM (songd2021enoising), we reformulated the noise injection method during sampling to ensure that at every timestep, the noise level of the latent variable remains consistent with the scheduler.
@@
To summarize, our main contributions are as follows:
  • We identify an issue of significant noise in images sampled via Flow-SDE (Figure 1). Through analysis, we introduce the concept of Coefficient-Preserving Sampling and prove that the original Flow-SDE fails to satisfy this requirement.
  • We propose a novel sampling formulation that adheres to the Coefficient-Preserving property. By generalizing DDIM to Flow Matching, the proposed algorithm generates high-fidelity images even under high noise levels.
  • We analyze the root cause of the excessive noise in Flow-SDE sampling, tracing it back to the Taylor expansion used in its derivation. We showed that this expansion not only introduces approximation errors but also induces numerical instability due to the inclusion of a $1/t$ term.
  • We experimentally verify that our method significantly facilitates both reward estimation and optimization, yielding results that substantially outperform those based on Flow-SDE sampling.
@@ 2.1 Alignment For Large Language Models
With the advent of Large Language Models (LLMs)(brown2020language; achiam2023gpt), Reinforcement Learning (RL) has garnered renewed attention. The Reinforcement Learning from Human Feedback (RLHF) framework(ouyang2022training; gao2023scaling), for instance, trains a reward model using human preference data, which in turn fine-tunes the LLM to better align with human expectations. As an alternative to computationally intensive policy gradient methods, Direct Preference Optimization (DPO)(rafailov2023direct) provides a more streamlined approach that directly trains the model on human preference data. More recently, advanced techniques have been applied to enhance multi-step reasoning. For example, OpenAI-o1 utilizes Proximal Policy Optimization (PPO)(schulman2017proximal) and DeepSeek-R1 employs Group Relative Policy Optimization (GRPO)(shao2024deepseekmath), both using verifiable rewards to improve the models’ capacity for extended reasoning via chain-of-thought.
@@ 2.2 Alignment for Diffusion
Analogous to autoregressive LLMs, both Diffusion(ho2020denoising; songd2021enoising; song2021score) and Flow Matching(lipman2022flow; liu2022flow) models usually construct their outputs via a multi-step sampling process. They can therefore be aligned using similar RL-based techniques. As such, diffusion models are compatible with optimization algorithms including DPO(wallace2024diffusion; dong2023raft; yang2024using), PPO-style policy gradients(black2023training; miao2024training; zhao2025score), and GRPO(liu2025flow; xue2025dancegrpo). However, unlike token-based models that involve discrete selection steps, the absence of quantization in the diffusion process permits an alternative training paradigm: the direct backpropagation of gradients through the full sampling trajectory xu2023imagereward.
@@ 2.3 Diffusion Sampler
Efficient sampling is a critical research area for diffusion models. Samplers can be broadly categorized by the numerical methods they adapt. Foundational approaches like DDPM (ho2020denoising) established the paradigm but were slow. Denoising Diffusion Implicit Models (songd2021enoising) provided one of the first major speed improvements by formulating a deterministic sampling process. Subsequently, a significant body of work has focused on applying and adapting sophisticated ordinary differential equation (ODE) solvers. For instance, first-order methods like the Euler solver (song2021score) offer speed at the cost of accuracy, while second-order methods like Heun’s method (karras2022elucidating) provide a better balance. High-order solvers, such as DPM-Solver (lu2022dpm; lu2025dpm), have become popular for their dramatic reduction in required sampling steps. Concurrent work zheng2025rethinking unifies all previous samplers by coefficient matrices, which are formulated by a similar rule with our proposed Coefficient-Preserving Sampling.
@@ 3 Preliminaries
Flow Matching Assume that $\bm{x}_{0}\sim X_{0}$ is sampled from the data distribution and $\bm{x}_{1}\sim X_{1}$ is a gaussian noise sample, Rectified Flow (liu2022flow) interpolates noised sample $\bm{x}_{t}$ as,
@@
where $\hat{\bm{v}}_{\theta}(\bm{x}_{t},t)$ is the estimated velocity. The hat ($\hat{\cdot}$) denotes that the value is model predicted in the following article.
Flow-GRPO and Dance-GRPO Reinforcement Learning relies on stochastic sampling to generate diverse samples. Flow-GRPO (liu2025flow) and Dance-GRPO (xue2025dancegrpo) introduce randomness into Flow Matching by converting the deterministic Flow-ODE into Flow-SDE,
@@
$A_{t}^{i}=\frac{R(\bm{x}_{0}^{i})-\text{mean}(\{R(\bm{x}^{i}_{0})\}_{i=1}^{G})}{\text{std}(\{R(\bm{x}_{0}^{i})\}_{i=1}^{G})}.$ (4)
Then GRPO (shao2024deepseekmath) optimizes the policy model by maximizing the following objective,
$\mathcal{L}(\theta)=\mathbb{E}_{\bm{x}^{i}\sim\pi_{\theta_{\text{old}}}}\frac{1}{G}\sum\limits_{i=1}^{G}\frac{1}{T}\sum\limits_{t=0}^{T-1}\Bigg(\mathop{min}\Big(r_{t}^{i}(\theta)A_{t}^{i},\text{clip}(r_{t}^{i}(\theta),1-\epsilon,1+\epsilon)A_{t}^{i})\Big)-\beta D_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\Bigg),$ (5)
where $r_{t}^{i}(\theta)=\frac{p_{\theta}(\bm{x}^{i}_{t-1}|\bm{x}_{t}^{i})}{p_{\theta_{\text{old}}}(\bm{x}^{i}_{t-1}|\bm{x}_{t}^{i})}$ and the KL loss term is defined as a closed form:
$D_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})=\frac{\|\overline{\bm{x}}_{t-\Delta t,\theta}-\overline{\bm{x}}_{t-\Delta t,\text{ref}}\|^{2}}{2\sigma_{t}^{2}\Delta t},$ (6)
where $\overline{\bm{x}}$ denotes the mean of predicted $\bm{x}$, which is implemented by removing the injected noise.
@@
where $\epsilon_{t}\sim{\mathcal{N}}({\bm{0}},{\bm{I}})$ is standard Gaussian noise independent of $\epsilon_{\theta}^{(t)}(\bm{x}_{t})$. For other notations, please refer to the DDIM paper (songd2021enoising). When $\sigma_{t}=\sqrt{(1-\alpha_{t-1})/(1-\alpha_{t})}\sqrt{1-\alpha_{t}/\alpha_{t-1}}$, the forward process becomes Markovian, and the generative process becomes a DDPM. When $\sigma_{t}=0$, the resulting model becomes an implicit probabilistic model (DDIM). For other $\sigma_{t}$, we call it DDIM with stochasticity.
@@ 4.1 Coefficients-Preserving Sampling
Referring to Equation 8, we can rewrite the Flow-ODE sampling function as,
@@
    The total noise level, defined as the standard deviation of a single multivariate noise or the root sum square (RSS) of the standard deviations of multiple independent noises, must align with the scheduler for all timesteps.
The definition of the total noise level relies on two key assumptions. First, we assume that the predicted noise terms, denoted as $\epsilon_{\theta}$ in Equation 7 or $\hat{\bm{x}}_{1}$ in Equation 8, adhere to the properties of a standard Gaussian distribution, e.g., zero mean and unit variance. This is a standard assumption in diffusion sampling algorithms songd2021enoising; karras2022elucidating; lu2022dpm, given that they necessitate replacing ground truth variables with predicted estimates during inference, despite the gap between training and testing. Second, we assume that the predicted noise is statistically independent of the newly injected noise $\epsilon_{t}$. This independence holds by construction, as $\epsilon_{t}$ is explicitly sampled from a fresh, independent Gaussian distribution at each timestep.
DDIM sampling is Coefficients-Preserving: In Equation 7, there are two independent noise terms, whose coefficients are $\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}$ and $\sigma_{t}$, so the final noise level is their RSS $\sqrt{1-\alpha_{t-1}}$. The sample coefficient is $\sqrt{\alpha_{t-1}}$, so the squared sum of the two coefficients is 1. These two coefficients exactly match the DDIM scheduler, whatever $\sigma_{t}$ is. Thus, we say the sampling procedure of DDIM is Coefficients-Preserving Sampling.
@@ 4.2 Flow-SDE is not Coefficients-Preserving Sampling
Recall Equation 3 and rewrite it into the similar form of Equation 7,
@@
$=\left(1-(t-\Delta t)\right)\hat{\bm{x}}_{0}+(t-\Delta t-\frac{\sigma_{t}^{2}\Delta t}{2t})\hat{\bm{x}}_{1}+\sigma_{t}\sqrt{\Delta t}\bm{\epsilon}.$ (10)
@@
Figure 2: The ideal noise level $t$ and SDE noise level (Equation 11) for Flow-GRPO and Dance-GRPO with $1000$, $16$, and 44 sampling steps. Except for the numerical problem around $t=0$ and $t=1$, the error of the noise level increases as the sampling step decreases.
In Figure 2, we plot the total noise level for both Flow-GRPO and Dance-GRPO. As we can see, the noise level mismatch problem is severe for both of them. Moreover, because of the $\frac{\sigma_{t}^{2}\Delta t}{2t}$ term in Equation 10, the error around $t=0$ is large for Dance-GRPO ($\sigma_{t}=\eta$). For Flow-GRPO, $\sigma_{t}=\eta\sqrt{\frac{t}{1-t}}$, the noise level is inaccurate around $t=1$. The problem becomes even worse when the sampling step is low, e.g. $4$ steps for FLUX.1-schnell.
@@ 4.3 Our Solution
The main problem of Flow-SDE is that the reduced noise level $\frac{\sigma_{t}^{2}\Delta t}{2t}$ cannot match the newly added noise level $\sigma_{t}\sqrt{\Delta t}$. Noticing that DDIM also injects noise into the sampling procedure while preserving the noise level (Figure 3.b), we consider referring to DDIM sampling to solve the problem. Assume that the newly added noise has a variance of $\sigma_{t}^{2}$, the coefficient of predicted noise should be $\sqrt{(t-\Delta t)^{2}-\sigma_{t}^{2}}$ to meet the requirement of the second condition of CPS. In this way, the sampling formulation is,
@@
For the injected noise level $\sigma_{t}$, the maximum value is $t-\Delta t$, or the sqrtsqrt term would have a negative radicand. To avoid the negative radicand, we propose to set $\sigma_{t}=(t-\Delta t)\sin(\frac{\eta\pi}{2})$. Then the sampling formulation becomes,
@@
To train with GRPO, we also need $p_{\theta}(\bm{x}^{i}_{t-1}|\bm{x}_{t}^{i})$, which is defined as (liu2025flow),
$\log p_{\theta}(\bm{x}^{i}_{t-1}|\bm{x}_{t}^{i})=-\frac{\|\bm{x}_{t-\Delta t}-\mu_{\theta}(\bm{x}_{t},t)\|^{2}}{2\sigma_{t}^{2}}-\log\sigma_{t}-\log\sqrt{2\pi},$ (14)
where $\mu_{\theta}(\bm{x}_{t},t)=\left(1-(t-\Delta t)\right)\hat{\bm{x}}_{0}+(t-\Delta t)\cos(\frac{\eta\pi}{2})\hat{\bm{x}}_{1}$ in our case. For each step, the $\log\sigma_{t}-\log\sqrt{2\pi}$ is a constant value that cancels out in $r_{t}^{i}(\theta)=\frac{p_{\theta}(\bm{x}^{i}_{t-1}|\bm{x}_{t}^{i})}{p_{\theta_{\text{old}}}(\bm{x}^{i}_{t-1}|\bm{x}_{t}^{i})}$. Moreover, we removed the $\sigma_{t}$ in the denominator to avoid division by zero or very small values in the last timestep. Thus, our definition of log-probability is as simple as,
$\log p_{\theta}(\bm{x}^{i}_{t-\Delta t}|\bm{x}_{t}^{i})=-\|\bm{x}_{t-\Delta t}-\mu_{\theta}(\bm{x}_{t},t)\|^{2}.$ (15)
Analytically, the normalization term $2\sigma_{t}^{2}$ disproportionately emphasizes the optimization of later timesteps, which involve less stochasticity. Removing this term reallocates greater weight to the earlier timesteps, which typically exhibit higher diversity and is crucial to Reinforcement Learning.
Meanwhile, the denominator in the KL loss function (Equation 6) should also be removed:
$D_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})=\|\mu_{\theta}(\bm{x}_{t})-\mu_{ref}(\bm{x}_{t})\|^{2}.$ (16)
Figure 3: (a): DDIM deterministic sampling process. Note that $\epsilon$ is a random Gaussian noise, which is almost orthogonal to the sample $\bm{x}_{0}$. Since $\sqrt{\alpha_{t}}^{2}+\sqrt{\beta_{t}}^{2}=1$, the trajectory is part of a quarter-circle at each step. (b): DDIM sampling process with stochasticity (Equation 7). $\epsilon_{t}$ is also a random Gaussian noise, which is almost orthogonal to $\bm{x}_{0}$ and $\epsilon_{\theta}$. (c): Flow matching ODE Sampler. The trajectory is a straight line at each step. (d): Our proposed Coefficients-Preserving Sampling (Equation 13). These figures are from blog wang2024zhihu.